OpenStudy (anonymous):

I have another that's kickin my butt. cos(2t)=1/2

OpenStudy (amistre64):

when is cos(t) = 1/2 ??

OpenStudy (anonymous):

yup and 0<t<2pi

OpenStudy (amistre64):

do you know your basic trig angles and their values?

OpenStudy (anonymous):

tryin to get better at it

OpenStudy (amistre64):

trig is like learning the multiplication tables; just need to get it memorized :)

OpenStudy (amistre64):

the basic triangle for this is the 60-30-90; 60 has a cos of 1/2

OpenStudy (anonymous):

cos(2t)=1/2 was the equation

OpenStudy (amistre64):

60 is pi/3 radians; so the question is: 2t = 60 t = 30 and t = 360-30 t = pi/6 and t = 330 = 12pi/6-pi/6 = 11pi/6

OpenStudy (amistre64):

if you know that cos(60) = 1/2 the 2t = 60 t = 30....